# ML4H 2025 Template: Proceedings Track

### v2

@@ 3 Solution 1: Model–Card Matching with Early Termination
Cutoffs in practice The purpose of per-stage cutoffs is to steer the workflow toward a more accurate outcome: either terminate early when plausible or return the best-matching model card (model code). Early termination can occur at any of the three stages—returning None or Other at Stage 1, Normal at Stage 2, or None at Stage 3—otherwise the pipeline outputs a model card id. To choose the cutoffs, we experimented on a broad set of out-of-distribution (OOD) clinical images, defined as any image outside MedGemma’s training/validation/test sets, drawn from the datasets in Section 4 and from public, license-free images Wikimedia Foundation (2025). The resulting thresholds were a low cutoff at Stage 1 ($0.10$), a moderate cutoff at Stage 2 ($0.30$), and a very small cutoff at Stage 3 ($0.025$). Figure 2 illustrates how the Stage 2 and Stage 3 thresholds can jointly flip the final outcome on the same renal ultrasound image Hansen et al. (2015). In this example, Stage 2 is incorrect (the top-two predicted abnormalities are wrong). If Stage 2 forwards its first candidate to Stage 3 (bottom panel), an incorrect model card is selected given the cutoff threshold. By contrast, if Stage 2 instead forwards the second candidate (pancreatic cancer; top panel), the workflow abstains, returning None. Thus, the interplay between the Stage 2 candidate choice and the Stage 3 cutoff can flip the decision from a wrong selection to safe early termination. Figure 3 illustrates two fixed–cutoff cases. Bottom: the very small Stage 3 cutoff allows the pipeline to select the second-ranked answer—a plausible model-card id for a chest X-ray with cardiomegaly from the NIH Long-Tail CXR dataset Wang et al. (2017)—rather than terminate and return None. This example shows that the Stage 3 threshold was calibrated so that, when sufficiently likely, a second candidate can be prioritized over the first (abstention) option. Top: when the pipeline is highly confident at each stage—for a histopathology image with carcinoma from Cruz-Roa et al. (2014)—the cutoffs are effectively inert and do not alter the outcome.
@@ 5 Discussion
MedGemma is pretrained on a fixed set of modalities and tasks; accordingly, it may not generalize to imaging types or findings outside its training distribution. In short, its inference and selection decisions are tied to its training data: if a modality or pathology is unseen or underrepresented, the model may fail to recognize it and either choose an imperfect model or default to “Other/None.” This risk can be mitigated by enlarging the pretraining corpus or by adopting continual-learning methods that add new domains while preserving prior knowledge (e.g., Li and Hoiem (2017)). Another limitation is that our current model card database supports up to two tasks per model (e.g., detection followed by classification); in reality, a case might benefit from multiple models (e.g., detection, classification and grading). Extending the pipeline to return a composite plan (multiple models in sequence) is an interesting direction. It is also worth noting that while our system outputs justifications for model selection, the interpretability of the VLM’s internal reasoning remains limited (a common issue with large neural models). Techniques like chain-of-thought prompting or explicability constraints could be applied so the model provides brief, stagewise reasoning—further improving transparency Wei et al. (2022). Another consideration is the choice of thresholds $\tau_{1},\tau_{2},\tau_{3}$. We converged to approximate values that work well based on our prompting strategy at each stage, but in practice they may need adjustment for different clinical contexts. A high-sensitivity application (where missing a disease is very costly) might use a lower $\tau_{2}$ to rarely terminate at “Normal”unless almost certain, whereas a high-specificity scenario might use a higher $\tau$ to avoid false alarms. Finally, our pipeline is across three stages because our repository lookup is primarily keyed on modality and one primary finding. As we incorporate more complex model cards (e.g., some models might handle combinations of findings or require patient metadata), the prompt design might need to extend to additional stages or a more free-form query. Fortunately, the modular nature of our approach means we can extend the prompting strategy as needed, and the answer selector concept can generalize to any number of stages by always examining the confidence of alternate responses. We outline two complementary directions for deployment: (1) automate fine-tuning end to end—so deployment proceeds with minimal data-scientist intervention—by automatically selecting the dataset linked to the model card chosen in solution 1 and automatically generating the prompts required for fine-tuning; and (2) reduce inference latency by applying pruning methods tailored to large language models (e.g., SparseGPT Frantar and Alistarh (2023)).

### v3

@@ 3 Solution 1: Model–Card Matching with Early Termination
Cutoffs in practice The purpose of per-stage cutoffs is to steer the workflow toward a more accurate outcome: either terminate early when plausible or return the best-matching model card (model code). Early termination can occur at any of the three stages—returning None or Other at Stage 1, Normal at Stage 2, or None at Stage 3—otherwise the pipeline outputs a model card id. To choose the cutoffs, we experimented on a broad set of out-of-distribution (OOD) clinical images, defined as any image outside MedGemma’s training/validation/test sets, drawn from the datasets in Section 4 and from public, license-free images Wikimedia Foundation (2025). The resulting thresholds were a low cutoff at Stage 1 ($0.10$), a moderate cutoff at Stage 2 ($0.30$), and a very small cutoff at Stage 3 ($0.025$). Figure 2 illustrates how the Stage 2 and Stage 3 thresholds can jointly flip the final outcome on the same renal ultrasound image Hansen et al. (2015). In this example, Stage 2 is incorrect (the top-two predicted abnormalities are wrong). If Stage 2 forwards its second candidate to Stage 3 (bottom panel), an incorrect model card is selected given the cutoff threshold. By contrast, if Stage 2 instead forwards the first candidate (liver cancer; top panel), the workflow abstains, returning None. Thus, the interplay between the Stage 2 candidate choice and the Stage 3 cutoff can flip the decision from a wrong selection to safe early termination. Figure 3 illustrates two fixed–cutoff cases. Bottom: the very small Stage 3 cutoff allows the pipeline to select the second-ranked answer—a plausible model-card id for a chest X-ray with cardiomegaly from the NIH Long-Tail CXR dataset Wang et al. (2017)—rather than terminate and return None. This example shows that the Stage 3 threshold was calibrated so that, when sufficiently likely, a second candidate can be prioritized over the first (abstention) option. Top: when the pipeline is highly confident at each stage—for a histopathology image with carcinoma from Cruz-Roa et al. (2014)—the cutoffs are effectively inert and do not alter the outcome.
@@ 5 Discussion
MedGemma is pretrained on a fixed set of modalities and tasks; accordingly, it may not generalize to imaging types or findings outside its training distribution. In short, its inference and selection decisions are tied to its training data: if a modality or pathology is unseen or underrepresented, the model may fail to recognize it and either choose an imperfect model or default to “Other/None.” This risk can be mitigated by enlarging the pretraining corpus or by adopting continual-learning methods that add new domains while preserving prior knowledge (e.g., Li and Hoiem (2017)). Another limitation is that our current model card database supports up to two tasks per model (e.g., detection followed by classification); in reality, a case might benefit from multiple models (e.g., detection, classification and grading). Extending the pipeline to return a composite plan (multiple models in sequence) is an interesting direction. It is also worth noting that while our system outputs justifications for model selection, the interpretability of the VLM’s internal reasoning remains limited (a common issue with large neural models). Techniques like chain-of-thought prompting or explicability constraints could be applied so the model provides brief, stagewise reasoning—further improving transparency Wei et al. (2022). Another design choice concerns the selection of thresholds $\tau_{1},\tau_{2},\tau_{3}$. We converged on approximate values that performed well given (i) our stage-specific prompting strategy; (ii) the diversity and structure of the model-card repository defined in Section 3; and (iii) the workflow’s subjective preference for the true final output—either None or a model-card ID—across different input images. In practice, these thresholds may require adjustment across clinical contexts. For high-accuracy applications, using the same CT of cardiomegaly in Figure 3, a workflow designer might prefer None over the closest model-card match (MODEL_13). Finally, our pipeline is across three stages because our repository lookup is primarily keyed on modality and one primary finding. As we incorporate more complex model cards (e.g., some models might handle combinations of findings or require patient metadata), the prompt design might need to extend to additional stages or a more free-form query. Fortunately, the modular nature of our approach means we can extend the prompting strategy as needed, and the answer selector concept can generalize to any number of stages by always examining the confidence of alternate responses. We outline two complementary directions for deployment: (1) automate fine-tuning end to end—so deployment proceeds with minimal data-scientist intervention—by automatically selecting the dataset linked to the model card chosen in solution 1 and automatically generating the prompts required for fine-tuning; and (2) reduce inference latency by applying pruning methods tailored to large language models (e.g., SparseGPT Frantar and Alistarh (2023)).